Kofi electronics bought a shipment of TV at a net price of $477.36 each, after discount of 15%, 10% and 4%. What is the list price?


1
Expert's answer
2021-03-08T13:59:39-0500

last price"=95/100\times 477.36\times 90/100\times 96/100=\$391.82"